OpenStudy (anonymous):

Find the solutions of the equation that are in the interval [0, 2π) tan2 x sin x = 3 sin x

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \tan^2x\sin x =3\sin x }\) link this ?

OpenStudy (solomonzelman):

*like* this?

OpenStudy (anonymous):

@SolomonZelman Yes! \[(\tan ^{2} x)(\sin x) = 3 \sin x\]

OpenStudy (solomonzelman):

yes, same thing:D

OpenStudy (anonymous):

I'm not sure what to do with the tan out front...

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \tan^2x\sin x =3\sin x }\) \(\large\color{black}{ \displaystyle \tan^2x\sin x -3\sin x =0}\) \(\large\color{black}{ \displaystyle \left(\tan^2x -3\right)\sin x =0}\)

OpenStudy (solomonzelman):

and the zero product property

OpenStudy (anonymous):

@SolomonZelman That's how far I got before I got stuck.

OpenStudy (anonymous):

@SolomonZelman So is it \[(\tan ^{2}x -3) = 0\] and \[\sin x = 0\] ?

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

Awesome thank you!

OpenStudy (solomonzelman):

yw.